Circular No. 7489

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

Z ANDROMEDAE J. A. Mattei, AAVSO, reports that this star has gone into outburst for

the first time since 1997 (cf. IAUC 6679), when the star peaked at m_{v} ∼ 9.7 on June 4 and began a slow fade later that month towards minimum at m_{v} ∼ 10.9 in 1998 March. A. Baransky, Kiev, Ukraine, has also reported the current outburst independently to the Bureau. Selected visual magnitude estimates (mostly from the AAVSO): Aug. 28.85 UT, 10.8 (Baransky); Sept. 1.10, 10.4 (M. Komorous, London, ON); 1.93, 10.8 (J. Ripero Osorio, Madrid, Spain); 2.81, 10.5 (Baransky); 3.24, 10.3 (E. Van Ballegoij, Oranjestad, Aruba); 4.09, 10.4 (B. Granslo, Oslo, Norway); 4.86, 10.1 (E. Muyllaert, Oostende, Belgium); 5.06, 10.1 (P. Dombrowski, Glastonbury, CT); 5.80, 10.0 (Baransky) 6.08, 10.1 (P. McDonald, Toronto, ON); 6.23, 9.8 (M. Simonsen, Macomb, MI); 7.07, 9.8 (J. Bortle, Stormville, NY); 7.92, 10.1 (Ripero Osorio); 8.10, 9.6 (Dombrowski).

V4642 SAGITTARII D. K. Lynch, R. J. Rudy, S. Mazuk, and C. Venturini, The Aerospace

Corporation; R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego; and T. Armstrong, Thomas K. Armstrong Development Corporation, McKinleyville, CA, report 0.8–2.5-μm spectrophotometry of V4642 Sgr (= N Sgr 2000) using the Lick Observatory 3-m telescope and the Aerospace Near Infrared Imaging Spectrograph (NIRIS) on July 20.31 UT, ∼ 161 days after peak brightness: “Although the fluorescently excited lines of O I persist, He II lines are now strong and coronal lines are now present. The latter include [S VIII] at 0.9913 μm, [Si VI] at 1.9629 μm, [Ca VIII] at 2.3214 μm, [Si VII] at 2.4827 μm, and possibly [Al IX] at 2.0469 μm. As is frequently the case, the He I line at 1.0830 μm is the strongest feature. There was no evidence of thermal emission from dust in the region 1–2.5 μm, although a significant, relatively flat continuum was present. J, H, and K magnitudes determined from the spectrum are 12.5, 12.7, and 11.6, respectively. The spectrum generally resembled that of V4633 Sgr (N Sgr 1998), as reported by Rudy et al. in 1999 (IAUC 7259).”

Additional visual magnitude estimates by A. Pearce, Nedlands, W. Australia (cf. IAUC 7381): Mar. 23.841 UT, 12.6; 27.844, 13.6; Apr. 1.836, 13.2; 9.831, 13.7; May 22.858, 14.7; June 1.842, 15.1; 11.837, 15.1.

2000 September 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green